OpenStudy (anonymous):

Please Help ill give medals and fans Use a graph and a table to solve the system. Check your answer. Enter your answer as the following example: (5, -6)

OpenStudy (anonymous):

\[ x - y = -1 & 4x - 2y = 2\]

OpenStudy (anonymous):

\[y - x = 0 & 8x + 4y = -24\]

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

so are you trying to find where the 2 lines intersect...?

OpenStudy (anonymous):

i have to find the point . Use a graph and a table to solve the system. Check your answer. Enter your answer as the following example: (5, -6) @campbell_st

OpenStudy (campbell_st):

ok... so create a table of values for each line by nominating solve x values then sibstituting them into the equation to find y x - y = -1 which can be written as y = x + 1 x: 0 : 1 : 2 ----------- y: 1 : 2 : 3 plot the points (0, 1), (1, 2) and (2, 3) draw a line through them next line 4x - 2y = 2 which can be written as y = 2x - 1 x: 0 : 1 : 2 ------------- y: -1 : 1: : 3 so plot the points (0, -1) (1, 1) and then (2, 3) draw a line through them. then find the point of intersection this site will allow you to graph the lines and identify the point of intersection https://www.desmos.com/calculator hope it helps